# TorchTitan: One-stop PyTorch native solution for production ready LLM pre-training

### v1

@@ 1 Introduction
LLM training challenges are being tackled from all sides. Training large language models (LLMs) at scale is a daunting task that requires a delicate balance of parallelism, computation, and communication, all while navigating intricate memory and computation tradeoffs. The massive resources required for training make it prone to GPU failures, underscoring the need for efficient recovery mechanisms and checkpointing strategies to minimize downtime (Eisenman et al., 2022; Wang et al., 2023; Gupta et al., 2024; Maurya et al., 2024; Wan et al., 2024). To optimize resource utilization and achieve elastic scalability, it is crucial to combine multiple parallelism techniques, including Data Parallel (Li et al., 2020; Rajbhandari et al., 2020; Zhang et al., 2022; Zhao et al., 2023), Tensor Parallel (Narayanan et al., 2021; Wang et al., 2022; Korthikanti et al., 2023), Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023; Fang and Zhao, 2024), and Pipeline Parallel (Huang et al., 2019; Narayanan et al., 2019, 2021; Tang et al., 2024). By stacking these parallelisms with memory and computation optimization techniques, such as activation recomputation (Chen et al., 2016; Korthikanti et al., 2023; He and Yu, 2023), mixed precision training (Micikevicius et al., 2018, 2022), and deep learning compilers (Bradbury et al., 2018; Yu et al., 2023; Li et al., 2024; Ansel et al., 2024), it is possible to maximize hardware utilization.
@@ 2.1.4 Pipeline Parallel
Pipeline Parallel (PP) views the model as a sequence of operations, chunking the operations (and the parameters used by them) into S stages which run on separate groups of devices. In the typical case, one stage represents a single model layer or a group of N adjacent model layers, but in theory it could be even be a partial layer. For the forward pass, a stage receives input activations (except stage 0), performs local computation, and sends output activations (except stage S - 1). The last stage performs a loss computation, and begins the backward pass, sending gradients in the reverse order through the pipeline. To improve efficiency, the input batch is broken into microbatches and a pipeline schedule overlaps computation on one microbatch with communication for others. TorchTitan enables a variety of pipeline schedules, with their schedules previously described in other works (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021; Tang et al., 2024).
@@ 4.1 Ongoing work: 4D parallelism and zero-bubble pipeline schedules
TorchTitan’s modular and extensible architecture enables the seamless integration of new techniques and optimizations. For instance, ongoing work includes incorporating Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023) to enable 4D parallelism, and leveraging the torch.distributed.pipelining package to support zero-bubble schedules (Tang et al., 2024). This demonstrates TorchTitan’s ability to adapt to evolving machine learning landscapes.

### v2

@@ 1 Introduction
LLM training challenges are being tackled from all sides. Training large language models (LLMs) at scale is a daunting task that requires a delicate balance of parallelism, computation, and communication, all while navigating intricate memory and computation tradeoffs. The massive resources required for training make it prone to GPU failures, underscoring the need for efficient recovery mechanisms and checkpointing strategies to minimize downtime (Eisenman et al., 2022; Wang et al., 2023; Gupta et al., 2024; Maurya et al., 2024; Wan et al., 2024). To optimize resource utilization and achieve elastic scalability, it is crucial to combine multiple parallelism techniques, including Data Parallel (Li et al., 2020; Rajbhandari et al., 2020; Zhang et al., 2022; Zhao et al., 2023), Tensor Parallel (Narayanan et al., 2021; Wang et al., 2022; Korthikanti et al., 2023), Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023; Fang and Zhao, 2024), and Pipeline Parallel (Huang et al., 2019; Narayanan et al., 2019, 2021; Qi et al., 2023). By stacking these parallelisms with memory and computation optimization techniques, such as activation recomputation (Chen et al., 2016; Korthikanti et al., 2023; He and Yu, 2023), mixed precision training (Micikevicius et al., 2018, 2022), and deep learning compilers (Bradbury et al., 2018; Yu et al., 2023; Li et al., 2024; Ansel et al., 2024), it is possible to maximize hardware utilization.
@@ 2.1.4 Pipeline Parallel
Pipeline Parallel (PP) views the model as a sequence of operations, chunking the operations (and the parameters used by them) into S stages which run on separate groups of devices. In the typical case, one stage represents a single model layer or a group of N adjacent model layers, but in theory it could be even be a partial layer. For the forward pass, a stage receives input activations (except stage 0), performs local computation, and sends output activations (except stage S - 1). The last stage performs a loss computation, and begins the backward pass, sending gradients in the reverse order through the pipeline. To improve efficiency, the input batch is broken into microbatches and a pipeline schedule overlaps computation on one microbatch with communication for others. TorchTitan enables a variety of pipeline schedules, with their schedules previously described in other works (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021; Qi et al., 2023).
@@ 4.1 Ongoing work: 4D parallelism and zero-bubble pipeline schedules
TorchTitan’s modular and extensible architecture enables the seamless integration of new techniques and optimizations. For instance, ongoing work includes incorporating Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023) to enable 4D parallelism, and leveraging the torch.distributed.pipelining package to support zero-bubble schedules (Qi et al., 2023). This demonstrates TorchTitan’s ability to adapt to evolving machine learning landscapes.